OpenStudy (anonymous):

2a a- b ----- + ------- a+b a addition and subtraction of rational expressions

OpenStudy (anonymous):

why dont you cross add... meaning that (a+b)+(a-b)=o and 2a+a= 3 a i not what you are trying to solve here but hope this helps

OpenStudy (anonymous):

we have to find LCD then rewrite rational expression with LCD then add then reduce

OpenStudy (anonymous):

ohh sorry im not in tht level yet :( im a freshman in honors :(

OpenStudy (anonymous):

oh lol thats okay... this is elementary college algebra... I just found this site so I am not sure what level it can really do... I'll just have to go in to school early tomorrow! Thank you though

OpenStudy (anonymous):

\[ \frac{2 a}{a+b}+\frac{a-b}{a} =\frac{3 a^2-b^2}{a (a+b)} \]

OpenStudy (anonymous):

\[ \frac{2 a}{a+b}+\frac{a-b}{a} =\frac{2a^2 + a^2 -b^2}{a (a+b)} =\frac{3 a^2-b^2}{a (a+b)} \]